What is the power dissipated in the 2.0 \Omega resistor in the circuit below?

A 10 W

B. 14.25 W

C. 12.5 W

D. 18.75 W
Answer #1

here we observed 3 ohm and 1 ohm are parallel connection

resistance R=3*1/3+1=0.75 ohm

resistance R ,4 and 2 ohm are series connection

the effective resistance Req=0.75+4+2=6.75 ohm

now we find the current on the battery

current i=18/6.75=2.67 A

now we find the power in 2 ohms

power P=2.67^2*2=14.25 W